Reports 1-1 of 1 Clear search Modify search
masayuki.nakano - 19:04 Wednesday 16 January 2019 (7671) Print this report
IMC recovery work and finesse measurement

[Ge, Nakano]

The IMC has been recovered. And the cavity pole was measured. I measured the TF from the servo output to the error point (output of mixer for the PDH). The AOM response is flat in this frequency region, so the TF should include only the TF of the cavity.

The result is shown in FIg.1. Note that because the optical gain changed, the DC gain is adjusted by my hand. The cavity pole seems to be slightly lower after the cleaning. 

However, it is still lower than the designed one. According to the fitting, the cavity pole is estimated as 8 kHz although the designed pole is 5.2 kHz. These correspond to the finesse of 320 and 520 and it seems to be too low. This might be because of the fitting error and I cannot believe it.

Images attached to this report
Search Help